OpenStudy (anonymous):

Multiplication Property of Negative One Prove that -1x X= -X a.−1×x+x=−1×x+1×x b.=(−1+1)(x) c.=0×x d. =0 e. =−x+x f.∴−1×x+x=−x+x g.∴−1×x=−x could anyone help me w/ the reasons for this??? like mult. prop. of = or sth...

OpenStudy (anonymous):

just say its common sense.

OpenStudy (anonymous):

Start from the idea that x=x Subtract x from both sides x-x=0 (by additive inverse property) subtract x from both sides again (x-x)-x= -x This gives 0-x=-x (again by additive inverse prop) factor out a -1 -1(0+x)=-x (by distributive law) collapse equation -1(x)=-x (by zero as neutral additive element)